OpenStudy (anonymous):

How to solve: 2log9 + 3log8

OpenStudy (anonymous):

Exact Form: log(41472) Decimal Form: 4.6177549

OpenStudy (anonymous):

ok, but how do get to your exact answer?

OpenStudy (anonymous):

I need to solve it without a calculator

OpenStudy (akashdeepdeb):

2 log 9 + 3 log 8 = log (9^2) + log (8^3) = log (81) + log (512) = log (41472) Let y = log (41472) 10^y = 41472 If you can approximate that; you have your answer. y ~ 4.617

OpenStudy (anonymous):

ah okay I see now, thanks alot